Solve the quadratic equations by taking the square roots: (X^2 - 5) + 4 = 0
Answer by ehollins1(8)

(x^2-5)+4=0
(x^2-5)+4=0
-4=0-4
x^2-5=-4
x^2-5=-4
+5=-4+5
x^2= 1
x= ±1
solution: x= ±1
*****check solution*****
x= ±1
x^2=1
(x^2-5)+4=0
(1^2-5)+4=0
(1-5)+4=0
-4+4=0
0=0